OpenStudy (anonymous):

Graph the function. Identify the vertex and axis of symmetry. f(x) = x2 +4x – 2 (1 point)

OpenStudy (anonymous):

how come when i do these i never get the one point?

OpenStudy (anonymous):

the first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{4}{2\times 1}=-2\]

OpenStudy (anonymous):

the second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(-2\)